Some ethylene glycol, HOCH2CH2OH, is added to your car’s cooling system along with 5.0


kg of water. If the freezing point of the water–glycol solution is −15.0 °C, what mass of


HOCH2CH2OH must have been added?


1
Expert's answer
2021-11-24T00:45:02-0500

∆Tf=Kf×m

0-(-15) =1.86×m

M=15/1.86

M=8.065